# WAXAL: A Large-Scale Multilingual African Language Speech Corpus

### v1

@@ Abstract
The advancement of speech technology has predominantly favored high-resource languages, creating a significant digital divide for speakers of most Sub-Saharan African languages. To address this gap, we introduce WAXAL, a large-scale, openly accessible speech dataset for 21 languages representing over 100 million speakers. The collection consists of two main components: an Automated Speech Recognition (ASR) dataset containing approximately 1,250 hours of transcribed, natural speech from a diverse range of speakers, and a Text-to-Speech (TTS) dataset with over 180 hours of high-quality, single-speaker recordings reading phonetically balanced scripts. This paper details our methodology for data collection, annotation, and quality control, which involved partnerships with four African academic and community organizations. We provide a detailed statistical overview of the dataset and discuss its potential limitations and ethical considerations. The WAXAL datasets are released at https://huggingface.co/datasets/google/WaxalNLP under the permissive CC-BY-4.0 license to catalyze research, enable the development of inclusive technologies, and serve as a vital resource for the digital preservation of these languages.
@@ 1 Introduction
Africa, home to 1.3 billion people and over 2,000 distinct languages, represents one of the most linguistically diverse regions on the planet. Despite this richness, Sub-Saharan Africa remains critically underserved by modern speech technology largely due to the scarcity of large-scale speech corpora in many of these languages. This data deficit is exacerbated by the unique linguistic characteristics of the region, such as tonal distinctions, complex morphology, and prevalent code-switching, which pose significant challenges for standard modeling paradigms. Without foundational datasets, it is impossible to train, evaluate, and adapt models to serve the educational, accessibility, and communication needs of these communities. To address this critical need, we present the WAXAL dataset, a new, large-scale resource for 21 Sub-Saharan African languages collected in partnership with four African academic and community partners and supported by Google’s technical mentorship. The collection features both ASR and TTS data for languages such as Acholi, Kiswahili, Luganda, Akan, and Ewe; ASR data for Shona, Fula, and Lingala; and TTS data for Hausa, Igbo, and Yoruba, among others. Our primary contributions are:
@@ 2 Related Works
Our work directly addresses these gaps by introducing a large-scale, unified dataset for 21 Sub-Saharan African languages. With approximately 1,250 hours of multi-speaker, transcribed speech for ASR and over 180 hours of high-quality, single-speaker recordings for TTS, our contribution significantly expands the data available for this linguistically diverse and technologically underserved region.
(a) ASR: Utterances per language.
(b) ASR: Hours of recordings per language
Fig. 2: WAXAL-ASR: Distribution of the ASR dataset sliced by language.
@@ 3 Data Collection Methodology
Fig. 3: WAXAL-ASR: Dist. of age and gender of speakers.
@@ 3.2 TTS Data Collection
(a) TTS: Recordings per language.
(b) TTS: Hours of recordings per language
Fig. 4: WAXAL-TTS: Distribution of the TTS dataset sliced by language.
@@ 4 Dataset Statistics
The WAXAL collection comprises distinct ASR and TTS datasets. The ASR dataset spans 14 languages, and the TTS dataset spans 10 languages. The total size of the released, transcribed ASR data is 1.7 TB, while the TTS data is 99 GB. Table 1 provides an overview.
Table 1: Statistics of the WAXAL ASR and TTS datasets.
@@
WAXAL-TTS 10 $\sim$186 17,660
5 Limitations and Ethical Considerations
@@ 6 Conclusion
The WAXAL dataset represents a significant step toward addressing the resource scarcity that has hindered the development of speech technologies for Sub-Saharan African languages. By providing over 1,400 hours of high-quality, annotated ASR and TTS data across 14 languages, we offer a foundational resource for building and evaluating models, conducting linguistic analysis, and fostering a more inclusive digital ecosystem. We have detailed our collaborative methodology, which prioritized local expertise and ethical data handling. We hope that this resource will empower researchers and developers to create practical, useful technologies that serve the communities these languages represent. The WAXAL datasets are being released publicly to the research community under the permissive CC-BY-4.0 license at https://huggingface.co/datasets/google/WaxalNLP.

### v3

# arXiv:2602.02734v3[eess.AS] 02 Mar 2026

@@ Abstract
The advancement of speech technology has predominantly favored high-resource languages, creating a significant digital divide for speakers of most Sub-Saharan African languages. To address this gap, we introduce WAXAL, a large-scale, openly accessible speech dataset for 24 languages representing over 100 million speakers. The collection consists of two main components: an Automated Speech Recognition (ASR) dataset containing approximately 1,250 hours of transcribed, natural speech from a diverse range of speakers, and a Text-to-Speech (TTS) dataset with over 235 hours of high-quality, single-speaker recordings reading phonetically balanced scripts. This paper details our methodology for data collection, annotation, and quality control, which involved partnerships with four African academic and community organizations. We provide a detailed statistical overview of the dataset and discuss its potential limitations and ethical considerations. The WAXAL datasets are released at https://huggingface.co/datasets/google/WaxalNLP under the permissive CC-BY-4.0 license to catalyze research, enable the development of inclusive technologies, and serve as a vital resource for the digital preservation of these languages.
@@ 1 Introduction
Africa, home to 1.3 billion people and over 2,000 distinct languages, represents one of the most linguistically diverse regions on the planet. Despite this richness, Sub-Saharan Africa remains critically underserved by modern speech technology largely due to the scarcity of large-scale speech corpora in many of these languages. This data deficit is exacerbated by the unique linguistic characteristics of the region, such as tonal distinctions, complex morphology, and prevalent code-switching, which pose significant challenges for standard modeling paradigms. Without foundational datasets, it is impossible to train, evaluate, and adapt models to serve the educational, accessibility, and communication needs of these communities. To address this critical need, we present the WAXAL dataset, a new, large-scale resource for 24 Sub-Saharan African languages collected in partnership with four African academic and community partners and supported by Google’s technical mentorship. The collection features both ASR and TTS data for languages such as Acholi, Kiswahili, Luganda, Akan, and Ewe; ASR data for Shona, Fula, and Lingala; and TTS data for Hausa, Igbo, and Yoruba, among others. Our primary contributions are:
@@ 2 Related Works
Our work directly addresses these gaps by introducing a large-scale, unified dataset for 24 Sub-Saharan African languages. With approximately 1,250 hours of multi-speaker, transcribed speech for ASR and over 180 hours of high-quality, single-speaker recordings for TTS, our contribution significantly expands the data available for this linguistically diverse and technologically underserved region.
(a) WAXAL-ASR: Transcribed hours per language.
(b) WAXAL-TTS: TTS hours of recordings per language
Fig. 2: WAXAL: Distribution of the Transcribed ASR and TTS datasets in hours sliced by language.
@@ 3 Data Collection Methodology
(a) ASR: #Utterances.
(b) ASR: Total hours of recordings.
Fig. 3: WAXAL-ASR: Total utterances and hours of recordings per language.
(a) ASR: Age distribution.
(b) ASR: Gender distribution
Fig. 4: WAXAL-ASR Demographics: Age and gender distribution of speakers.
@@ 4 Dataset Statistics
The WAXAL collection comprises distinct ASR and TTS datasets. The ASR dataset spans 14 languages, and the TTS dataset spans 13 languages. The total size of the released, ASR data is 1.7 TB, while the TTS data is 99 GB. Table 1 provides an overview.
Table 1: Statistics of the Transcribed ASR and TTS datasets.
@@
WAXAL-TTS 13 $\sim$235 $\sim$17,660
5 Limitations and Considerations
@@ 6 Conclusion
The WAXAL dataset represents a significant step toward addressing the resource scarcity that has hindered the development of speech technologies for Sub-Saharan African languages. By providing around 1,500 hours of high-quality, annotated ASR and TTS data across 24 languages, we offer a foundational resource for building and evaluating models, conducting linguistic analysis, and fostering a more inclusive digital ecosystem. We have detailed our collaborative methodology, which prioritized local expertise and ethical data handling. We hope that this resource will empower researchers and developers to create practical, useful technologies that serve the communities these languages represent. The WAXAL datasets are being released publicly to the research community under the permissive CC-BY-4.0 license at https://huggingface.co/datasets/google/WaxalNLP.